# Self-Managing DRAM: A Low-Cost Framework for Enabling Autonomous and Efficient in-DRAM Operations

### v5

@@ Abstract
The memory controller is in charge of managing DRAM maintenance operations (e.g., refresh, RowHammer protection, memory scrubbing) in current DRAM chips. Implementing new maintenance operations often necessitates modifications in the DRAM interface, memory controller, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, leading to slow progress in DRAM systems.
In this paper, our goal is to 1) ease the process of enabling new DRAM maintenance operations and 2) enable more efficient in-DRAM maintenance operations. Our idea is to set the memory controller free from managing DRAM maintenance. To this end, we propose Self-Managing DRAM (SMD), a new low-cost DRAM architecture that enables implementing new in-DRAM maintenance mechanisms (or modifying old ones) with no further changes in the DRAM interface, memory controller, or other system components. We use SMD to implement new in-DRAM maintenance mechanisms for three use cases: 1) periodic refresh, 2) RowHammer protection, and 3) memory scrubbing. Our evaluations show that, SMD-based maintenance operations significantly improve the system performance and energy efficiency while providing higher reliability compared to conventional DDR4 DRAM. On average across all workloads, a combination of SMD-based maintenance mechanisms that perform refresh, RowHammer protection, and memory scrubbing achieve 9.2% speedup and consume 6.2% less DRAM energy. We make SMD source code openly and freely available at [133].
@@ 1 Introduction
DRAM is the prevalent main memory technology used in almost all computer systems. Advances in manufacturing technology and DRAM design enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [78, 101, 105, 108]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrink, ensuring reliable and efficient DRAM operation becomes an even more critical challenge[4, 13, 44, 56, 62, 70, 83, 91, 101, 103, 108, 115, 124]. A modern DRAM chip typically requires three types of maintenance operations for reliable and secure operation: 1) DRAM refresh, 2) RowHammer protection, and 3) memory scrubbing.111Memory scrubbing is not always required in consumer systems but often used by cloud service providers [47, 98, 132, 94, 131, 141, 126, 35].
DRAM Refresh. Due to the short retention time of DRAM cells, all DRAM cells need to be periodically refreshed to ensure reliable operation. In existing DRAM chips, a cell is typically refreshed every 64 or below [96, 50, 51] and every 32 or above [52, 53]. A DRAM chip consists of billions of cells, and therefore refreshing such a large number of cells incurs large performance and energy overheads [86, 85, 124]. These overheads are expected to increase as DRAM chips become denser [16, 56, 86]. Many ideas have been proposed to mitigate the DRAM refresh overhead [16, 86, 124, 107, 6, 8, 23, 26, 34, 46, 55, 63, 90, 64, 87, 99, 106, 117, 145, 57, 58, 59, 150, 121, 125, 39, 68, 110, 74].
RowHammer Protection. Small and densely-packed DRAM cells suffer from the RowHammer vulnerability [70], which leads to disturbance failures that stem from electro-magnetic interference created by repeatedly activating (i.e., hammering) a DRAM row (i.e., aggressor row) many times [158, 33, 54, 113, 114, 127, 151, 157, 112, 102, 154]. Such rapid row activations lead to bit flips in DRAM rows that are physically nearby an aggressor row. Many works propose mechanisms to protect DRAM chips against RowHammer [3, 70, 5, 66, 81, 116, 139, 143, 146, 156, 159, 155, 36, 39, 129, 73, 12, 149, 92, 67, 28].
Memory Scrubbing. System designers typically use Error Correcting Codes (ECC) [38, 25] to improve system reliability. Moreover, DRAM vendors recently started to equip DRAM chips with on-die ECC to improve factory yield [53, 120, 119, 118, 109]. Detection and correction capability of an ECC scheme is limited. For example, Single-Error-Correction Double-Error-Detection (SECDED) ECC detects up to two bit errors and corrects one bit error in a codeword. To prevent independent single-bit errors from accumulating into multi-bit errors, memory scrubbing [47, 98, 132, 94, 131, 141, 126, 35] is commonly employed: the entire memory is scanned to correct single-bit errors.
We consider periodic refresh, RowHammer protection, and memory scrubbing as DRAM maintenance operations needed for reliable and secure operation of modern and future DRAM chips. An ideal maintenance operation should efficiently improve DRAM reliability and security with minimal system performance and energy consumption overheads.
As DRAM technology node scales to smaller sizes, the reliability and security of DRAM chips worsen. As a result, new DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [53, 52, 3]) and introducing new types of maintenance operations (e.g., targeted refresh [41, 31, 48] and DDR5 RFM [53] as RowHammer defenses). As the intensity and number of different maintenance operations increase, it becomes increasingly more challenging to keep the overhead of the maintenance operations low.
@@
Unfortunately, modifying the DRAM maintenance operations is difficult due to the current rigid DRAM interface that places the Memory Controller (MC) completely in charge of DRAM control. Implementing new maintenance operations (or changing existing ones) often necessitate modifications in the DRAM interface, MC, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, leading to slow progress in DRAM systems. For example, there was a five-year gap between the DDR3 [49] and DDR4 [50] standards, and eight-year gap between the DDR4 [50] and DDR5 [53] standards. In addition to waiting for a new standard to come out, DRAM vendors also need to push their proposal regarding the maintenance operations through the JEDEC committee so the new standard includes the desired changes to enable new maintenance operations. Thus, we believe a flexible DRAM interface that allows DRAM vendors to quickly implement custom in-DRAM maintenance mechanisms would help to develop more reliable and secure DRAM chips.
Our goal is to 1) ease the process of implementing new DRAM maintenance operations and 2) enable more efficient in-DRAM maintenance operations. We believe that enabling autonomous in-DRAM maintenance operations would encourage innovation, reduce time-to-market, and ease adoption of novel mechanisms that improve DRAM efficiency, reliability, and security. DRAM architects will quickly deploy their new in-DRAM solutions to tackle reliability and security problems of DRAM without demanding further changes in the DRAM interface, MC, or other system components. Autonomous in-DRAM maintenance operations are appealing to DRAM vendors because they would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to processing unit vendors or other DRAM vendors).
We propose Self-Managing DRAM (SMD), a new DRAM architecture that enables DRAM vendors to implement new maintenance operations and modify the existing ones without the need to make further changes in the DRAM interface, MC, or other system components. The key idea of SMD is to allow a DRAM chip to autonomously perform maintenance operations by preventing memory accesses to DRAM regions that are under maintenance. To perform a read or write access, the MC must first activate (i.e., open) a DRAM row. Thus, SMD successfully prevents accesses to a DRAM region (e.g., a specific DRAM bank) that is under maintenance by rejecting a row activation command (i.e., ACT) issued by the MC. After the DRAM chip completes the maintenance operation, the MC re-issues an ACT to the same row.
An SMD chip autonomously performs maintenance operations. For example, the MC never issues REF commands to an SMD chip. In fact, the MC does not even know which maintenance operations an SMD chip performs. In general, a maintenance mechanism in an SMD chip operates in three steps. First, the maintenance mechanism locks a DRAM region where it will perform maintenance. SMD prevents the MC from accessing the locked region by rejecting row activation commands (i.e., ACT) that target the locked region. Second, the maintenance mechanism performs a maintenance operation, i.e., refreshes rows within the region, performs RowHammer mitigation, or corrects bit flips via scrubbing. Third, after completing the maintenance, the maintenance mechanism releases the locked DRAM region to allow the MC to access it.
@@
To enable practical adoption of SMD, we implement it with low-cost modifications in existing DRAM chips and MCs. Although SMD requires simple changes in the MC to enable re-issuing a previously-rejected ACT command, SMD simplifies MC design as it is no longer responsible for managing DRAM maintenance operations. First, to inform the MC that a row activation is rejected, SMD adds a single uni-directional pin to a DRAM chip. Second, in a DRAM chip, SMD implements a Lock Controller for managing regions under maintenance and preventing access to them. In section 7.6, we show that the core components of SMD have low DRAM chip area overhead (1.63% of a DRAM chip) and latency cost (only 0.4% additional row activation latency).
We demonstrate the versatility and example benefits of SMD by implementing maintenance mechanisms for three use cases: 1) periodic refresh, 2) RowHammer protection, and 3) memory scrubbing.
First, we implement two maintenance mechanisms for DRAM refresh. Fixed-Rate Refresh (SMD-FR) uniformly refreshes the entire DRAM chip at a fixed refresh period. Compared to conventional DRAM refresh, our approach has the benefit of allowing accesses to DRAM regions that are not being refreshed. Variable Refresh (SMD-VR) refreshes DRAM rows with different periods based on the data retention capability of each row, further lowering the refresh overheads.
Second, we implement three maintenance mechanisms for RowHammer protection. Probabilistic RowHammer Protection (SMD-PRP) protects a DRAM chip against RowHammer by refreshing the neighbors of an activated row with a low probability. SMD-PRP+ uses area-efficient Bloom Filters to identify potential aggressor rows and refresh the neighbors of only such rows, eliminating unnecessary neighbor row refreshes when the activation count of an aggressor row is low. SMD-DRP is a deterministic RowHammer protection mechanism that keeps track of the most-frequently activated DRAM rows and refreshes their neighbors when activated more than a certain threshold. Existing DRAM chips use on-die Target Row Refresh (TRR) that piggybacks TRR-induced refreshes to regular refresh commands, which does not incur performance overhead but provides limited time for RowHammer protection [41, 48, 31]. Compared to TRR, SMD-PRP, SMD-PRP+, and SMD-DRP can be configured to efficiently refresh neighbor rows at the desired rate to strengthen the RowHammer protection.
Third, we implement Memory Scrubbing (SMD-MS), which scans the DRAM at a fixed time interval to correct single-bit errors. Compared to conventional memory scrubbing that is performed by the MC, SMD-MS has significantly lower scrubbing overhead, especially at low scrub periods (e.g., ).
We demonstrate the implementation of five example SMD maintenance mechanisms that resemble mechanisms implemented in existing DRAM chips. However, SMD chip designers have the freedom to implement different maintenance mechanisms without exposing any information to the MC or other system components.
Our evaluations show that a system equipped with SMD DRAM chips, which use combination of SMD-based maintenance mechanisms for refresh, RowHammer protection, and memory scrubbing, achieves 9.2% speedup and 6.2% lower DRAM energy on average compared to a typical DDR4 system, which performs only DRAM refresh, across all workloads. When using () refresh period, SMD-FR/SMD-VR provides 8.7%/9.2% (19.5%/20.7%) average speedup compared to conventional DDR4 refresh across 20 memory-intensive four-core workloads, achieving 86.8%/91.2% (86.6%/92.3%) of performance provided by a hypothetical DRAM that does not require any refresh.
@@
We expect and hope that SMD will inspire researchers to develop new DRAM maintenance mechanisms and enable practical adoption of innovative ideas by providing freedom to both DRAM and MC designers. We make SMD source code openly and freely available at [133].
@@
We use SMD to implement efficient DRAM maintenance mechanisms for three use cases: periodic refresh, RowHammer protection, and memory scrubbing.
We rigorously evaluate the performance and energy of the new maintenance mechanisms. SMD provides large performance and energy benefits while also improving reliability across a variety of systems and workloads.
@@ 2 Background
We provide a brief background on DRAM. For more detailed descriptions of DRAM organization and operation, we refer the reader to many prior works [15, 16, 17, 18, 39, 40, 70, 71, 76, 79, 77, 78, 80, 86, 85, 134, 135, 136, 137, 160, 62, 31, 89, 155, 152, 19, 120, 9, 162, 61, 60, 111, 72, 154, 11].
@@ 2.1 DRAM Organization
A typical computing system has one or more DRAM channels, where each channel has an independently operating I/O bus. As Fig. 1 illustrates, a Memory Controller (MC) interfaces with one or multiple DRAM ranks via the channel’s I/O bus. A DRAM rank consists of a group of DRAM chips that operate in lockstep. Because the I/O bus is shared across ranks, accesses to different ranks happen in serial manner. A DRAM chip is divided into multiple DRAM banks, each of which is further divided into multiple two dimensional DRAM cell arrays, called subarrays. Within each subarray, DRAM cells are organized as rows and columns. A DRAM cell connects to a sense amplifier via a bitline, and all sense amplifiers in the subarray form a row buffer. Within each DRAM cell, the data is stored as electrical charge on the capacitor and accessed through an access transistor. To perform a DRAM access, a row first must be activated (i.e., opened), which loads row’s data into the row buffer. A row activation is performed by enabling the wordline that corresponds to the row address that the MC provides along with the ACT command. To enable the corresponding wordline, the global row decoder first partially decodes the row address, and the local row decoder selects a wordline based on the partially decoded address provided by the global row decoder.
@@ 2.2 Accessing DRAM
To access a DRAM cell, the MC first needs to activate (i.e., open) the row that contains the cell by issuing an ACT command to a DRAM bank. A DRAM chip performs row activation in two steps. First, the wordline that corresponds to the row address is enabled to connect each cell capacitor in the row to its corresponding bitline. This causes charge sharing between a cell capacitor and a bitline. Due to the charge sharing, the bitline voltage slightly deviates from its pre-activation voltage level. Second, a sense amplifier detects the bitline voltage deviation and gradually restores the cell capacitor back to its original charge level. The data in the row buffer becomes accessible by a RD/WR command when the bitline and cell voltage is restored to a certain level. Therefore, the MC should follow an ACT with one or multiple RD/WR commands at least after the activation latency, which is typically denoted as tRCD.
To open a row in a bank that already has another row open, the MC first precharges the bank (i.e., closes the open row) by issuing a PRE command. The MC follows an ACT with a PRE at least after the restoration latency, which is typically denoted as tRAS. A DRAM chip performs a precharge operation in two steps. First, the DRAM chip disables the wordline of the activated row to disconnect its cells from the bitlines, and thus the row buffer. Second, the precharge circuitry, which is part of a sense amplifier, precharges the bitline to prepare it for the next row activation. The MC follows a PRE with an ACT to activate a new row at least after the precharge latency, which is typically denoted as tRP.
2.3 Periodic DRAM Refresh
A DRAM cell capacitor loses its charge over time [85, 86, 130], and losing too much charge results in a bit flip. The retention time of a cell refers to the duration for which its capacitor stores the correct value. Cells in a DRAM chip have different retention times, ranging from milliseconds to several hours [42, 56, 57, 65, 76, 84, 85, 86, 121, 124, 150]. However, to ensure data integrity, all DRAM cells are uniformly refreshed at fixed time intervals (typically 64, 32, or even ), called refresh window (tREFW). To perform periodic DRAM refresh, the MC issues a REF command at a fixed interval, denoted as refresh interval (tREFI) (typically 7.8 or ). A DRAM chip refreshes several (e.g., ) rows upon receiving a single REF command, and typically REF commands refresh the entire DRAM chip in a refresh window.

### v6

@@ Abstract
The memory controller is in charge of managing DRAM maintenance operations (e.g., refresh, RowHammer protection, memory scrubbing) to reliably operate modern DRAM chips. Implementing new maintenance operations often necessitates modifications in the DRAM interface, memory controller, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, likely leading to slow progress in the adoption of new architectural techniques in DRAM chips.
We propose a new low-cost DRAM architecture, Self-Managing DRAM (SMD), that enables autonomous in-DRAM maintenance operations by transferring the responsibility for controlling maintenance operations from the memory controller to the SMD-chip. To enable autonomous maintenance operations, we tweak the DRAM interface with a single, simple modification, such that an SMD chip rejects memory controller accesses to DRAM regions (e.g., a subarray or a bank) under maintenance, while allowing memory accesses to other DRAM regions. Thus, SMD enables 1) implementing new in-DRAM maintenance mechanisms (or modifying existing ones) with no further changes in the DRAM interface, memory controller, or other system components, and 2) overlapping the latency of a maintenance operation in one DRAM region with the latency of accessing data in another.
We evaluate SMD and show that it 1) can be implemented without modifications to the physical DDRx interface with low latency (0.4% of row activation latency) and area (1.6% of a DRAM chip) overhead, 2) achieves 4.1% average speedup across 20 four-core memory-intensive workloads over a DDR4-based system/DRAM co-design technique that intelligently parallelizes maintenance operations with memory accesses, and 3) guarantees forward progress for rejected memory accesses.
@@ 1 Introduction
Advances in manufacturing technology enable increasingly smaller DRAM cell sizes, continuously reducing cost per bit of a DRAM chip [1, 2, 3, 4]. However, as a DRAM cell becomes smaller and the distance between adjacent cells shrink, ensuring reliable and efficient DRAM operation becomes an even more critical challenge [5, 6, 7, 8, 9, 10, 11, 12, 2, 13, 4, 14, 15]. A modern DRAM chip requires three types of maintenance operations (described in detail in section 5) for reliable and secure operation: 1) DRAM refresh [16, 17, 15, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43], 2) RowHammer protection [44, 10, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 40, 56, 57, 58, 59, 60, 61, 62, 63], and 3) memory scrubbing.111Memory scrubbing is not always required in consumer systems but often used in cloud systems [64, 65, 66, 67, 68, 69, 70, 71]. New DRAM chip generations necessitate making existing maintenance operations more aggressive (e.g., lowering the refresh period [72, 73, 44]) and introducing new types of maintenance operations (e.g., targeted refresh [74, 75, 76] and DDR5 RFM [72] as RowHammer defenses).
Two problems likely hinder effective and efficient maintenance mechanisms from being adopted in modern and future DRAM-based computing systems. First, it is difficult to modify existing maintenance mechanisms and introduce new maintenance operations because doing so often necessitates changes to the DRAM interface, which takes a long time (due to various issues related to standardization and agreement across many vendors). Second, it is challenging to keep the overhead of DRAM maintenance mechanisms low as DRAM reliability characteristics worsen and DRAM chips require more aggressive maintenance operations. We expand on the two problems in the next two paragraphs.
@@
Implementing new maintenance operations (or changing existing ones) often necessitates modifications in the DRAM interface, MC, and potentially other system components. Such modifications are only possible with a new DRAM standard, which takes a long time to develop, likely leading to slow progress in the adoption of new architectural techniques in DRAM chips. For example, there was a five-year gap between the DDR3 [77] and DDR4 [78] standards, and eight-year gap between the DDR4 [78] and DDR5 [72] standards. While developing a new standard, DRAM vendors need to push their proposal regarding the maintenance operations through the JEDEC committee (which is the case even if the gap between subsequent DRAM standards were to substantially reduce) such that the new standard includes the desired changes to enable new maintenance operations. Thus, a flexible DRAM interface that decouples improvements to DRAM maintenance operations from interface changes (and hence the timeline needed to enable such changes) would allow DRAM designers to quickly implement custom in-DRAM maintenance mechanisms and develop more reliable and secure DRAM chips.
A maintenance operation triggers more frequently and takes longer with worsening DRAM reliability characteristics and increasing density. While a maintenance operation executes, DRAM is unavailable to serve memory requests (e.g., a periodic refresh operation in DDR4 prevents access to a DRAM rank for up to [78]). As a result, maintenance operations can significantly degrade system performance by delaying critical memory requests. Thus, enabling autonomous maintenance operations without delaying critical memory requests would improve system performance and energy efficiency.
Our goal is to 1) ease and accelerate the process of implementing new in-DRAM maintenance operations and 2) enable more efficient maintenance operations. To this end, we propose Self-Managing DRAM (SMD), a new DRAM architecture that enables implementing new DRAM maintenance operations and modifying the existing ones with a single, simple interface change that eliminates the need for future changes in the DRAM interface, the MC, or other system components.
SMD’s key idea is to allow a DRAM chip to autonomously and efficiently perform maintenance operations by preventing memory accesses only to a relatively small, under-maintenance DRAM region, i.e., a designated section in a DRAM chip (e.g., a DRAM subarray or a bank), while allowing memory accesses to other DRAM regions.222SMD has the key property that the MC does not even know which maintenance operations an SMD chip performs. This property of SMD could also be appealing to DRAM vendors because it would allow vendor-, and chip-generation-specific DRAM resilience characteristics (e.g., the degree and distribution of RowHammer vulnerability and retention failures) to remain undisclosed (i.e., not exposed to other DRAM vendors or system designers).
To prevent access to an under-maintenance DRAM region, an SMD chip rejects a row activation command (i.e., ACT) issued by the MC to the region. To ensure that a DRAM region cannot be locked for too long, SMD enforces a minimum delay (called ARI) between consecutive maintenance operations targeting the same region. We comprehensively study SMD and show that it ensures forward progress for memory requests (section 4.5).
While a DRAM region is under maintenance, the MC can access other regions. This way, a majority of memory accesses are not delayed by maintenance operations, as SMD overlaps the latency of a maintenance operation in a DRAM region with a memory accesses to another DRAM region.
@@
To enable practical adoption of SMD, we implement it with low-cost modifications in existing DRAM chips and MCs. First, to inform the MC that a row activation is rejected, SMD uses a single uni-directional pin on the DRAM chip (which already exists in DDR4/5, see section 7.4). Second, SMD implements a Lock Controller in a DRAM chip for managing regions under maintenance and preventing access to them. Third, SMD adds a new row address latch for every lock region in a DRAM chip to enable accessing one lock region while another is under maintenance (building on the basic design proposed in prior works [79, 80, 16]). In section 7.4, we show that these modifications have low DRAM chip area overhead (1.6% of a DRAM chip) and latency cost (only 0.4% additional row activation latency).
We demonstrate the practicality and versatility of SMD by implementing in-DRAM maintenance mechanisms for DRAM refresh (section 5.1), RowHammer protection (section 5.2), and memory scrubbing (section 5.3). We rigorously evaluate SMD’s impact on system performance and energy efficiency using cycle-accurate memory system simulations (using Ramulator [81, 82]), executing a diverse set of 62 single-core and 60 four-core workloads. We compare SMD-based implementations of the evaluated maintenance mechanisms to their MC-based implementations. We make three key observations from our evaluation. First, an SMD chip that implements lock regions at subarray granularity (i.e., a maintenance mechanism locks a small designated section in a bank) and implements all three maintenance mechanisms (refresh, RowHammer protection, and scrubbing) 1) improves average performance across all memory-intensive four-core workloads by 8.6% compared to a baseline system with a DDR4 chip that supports bank-level refresh and a memory controller that intelligently schedules maintenance operations to avoid delaying main memory accesses (DARP [16]) and 2) outperforms a system in which the DDR4 chip can concurrently perform a maintenance operation in a subarray and a memory access in another subarray (maintenance-access parallelization) and the memory controller can intelligently exploit maintenance-access parallelization (e.g., similar to DSARP’s [16] refresh-access parallelization) by 4.1%.
Second, SMD (with maintenance-access parallelization) reduces DRAM energy consumption by 4.3% on average as it eliminates DRAM commands issued by the memory controller to perform maintenance operations and reduces total execution time. Third, a very low chip area overhead (0.001% of a DRAM chip) implementation of SMD that locks regions at bank granularity (i.e., a maintenance mechanism locks a DRAM bank) and performs only refresh operations for maintenance (similar to DDR4) induces 4.5% average slowdown due to rejected row activations across all memory-intensive four-core workloads compared to the same baseline system (DARP [16]). We conclude that SMD practically provides better performance and energy benefits, without having to modify the DRAM interface (e.g., by adding new timing parameters to the standard) for every new type of maintenance mechanism, than system-DRAM co-design techniques that enable refresh-access parallelization (e.g., [16, 80]).
@@
We expect and hope that SMD will inspire researchers to develop new DRAM maintenance mechanisms that more efficiently tackle the reliability and security problems of DRAM and to enable practical adoption of innovative ideas by providing freedom to both DRAM and MC designers. We will publicly release all SMD source code and data so that others can replicate and build on the work.
@@
We use SMD to implement efficient DRAM maintenance mechanisms for three use cases: DRAM refresh, RowHammer protection, and memory scrubbing.
We rigorously evaluate the performance and energy of SMD-based maintenance mechanisms. SMD provides large performance and energy benefits while also improving reliability across a variety of systems and workloads.
@@ 2.1 DRAM Organization
A computing system has one or more DRAM channels, where each channel has an independently operating I/O bus. As Fig. 1 illustrates, a Memory Controller (MC) interfaces with one or multiple DRAM ranks via the channel’s I/O bus. A DRAM rank consists of a group of DRAM chips that operate in lockstep. Because the I/O bus is shared across ranks, accesses to different ranks happen in serial manner. A DRAM chip is divided into multiple DRAM banks, each of which is further divided into multiple two dimensional DRAM cell arrays, called subarrays. Within each subarray, DRAM cells are organized as rows and columns. A DRAM row consists of hundreds of DRAM cells which connect to the same wordline. A DRAM cell connects to a sense amplifier via a bitline, and all sense amplifiers in the subarray form a row buffer. Within each DRAM cell, the data is stored as electrical charge on the capacitor and accessed through an access transistor.
@@ 2.2 Accessing DRAM
A row must be activated (i.e., opened), to load the row’s data into the row buffer. A row activation is performed by enabling the wordline corresponding to the row address of the ACT command. To enable the corresponding wordline, the global row decoder first partially decodes the row address, and the local row decoder selects a wordline based on the partially decoded address provided by the global row decoder. Enabling the wordline copies the row’s data to the row buffer.
Data in the row buffer is accessible by a RD/WR. The MC follows an ACT with one or multiple RD/WR commands at least after the activation latency (tRCD). The MC issues a PRE command to close a row, after which it can activate a new row from the same bank. The time interval for issuing a PRE after an ACT must be at least restoration latency (tRAS). The MC follows a PRE with an ACT to activate a new row at least after the precharge latency (tRP).